during a hot air balloon ride the people were blown 50m north 625m west and50m south. what is the distance and the displacement of the balloon.
1
Expert's answer
2012-06-05T10:35:33-0400
distance is D = 50 + 625 + 50 = 725 m.
displacement is S = sqrt( (50 - 50)^2
+ (625)^2 ) = 625 m.